OpenStudy (missmeow):

Here is what happened. So I took this quiz, and bombed it. It really sucked. So I started to look over the questions, and it said this one is wrong (I chose C.) Which of these is a simplified form of the equation 7x + 4 = 5 + 3x + 1x? A. 3x = 9 B. 3x = 1 C. 11x = 9 D. 11x = 1 But the only problem is that I've got confirmation that C is in fact the right answer. So my question is, was I wrong?

OpenStudy (missmeow):

And if I was, how can I solve this equation?

OpenStudy (texaschic101):

7x + 4 = 5 + 3x + 1x 7x + 4 = 5 + 4x 7x - 4x = 5 - 4 3x = 1

OpenStudy (missmeow):

In this part, 7x - 4x = 5 - 4, how did the plus sign become a minus sign?

OpenStudy (texaschic101):

first, i subtracted 4x from both sides, and then i subtracted 4 from both sides ? do u understand ?

OpenStudy (missmeow):

Yes, thank you!

OpenStudy (texaschic101):

yw